Bank X has Sh. 100M of assets and Sh. 10M of Equity bank. Bank Y has Sh. 100M in assets and Sh. 4Min equity.
Calculate the Equity Multiplier of each bank. (5 Marks)

b) If the two banks have 1%, return on assets; calculate the return on equity of bank X and Y and comment.
1
Expert's answer
2018-10-31T13:53:09-0400

Bank X has Sh.100M of assets and Sh.10M of Equity. Bank Y has Sh.100M in assets and Sh.4M in equity.

a) Calculate the Equity Multiplier of each bank.

Equity multiplier = Total Asset/Total Equity

Equity multiplier X = 100/10 = 10,

Equity multiplier Y = 100/4 = 25.

b) If the two banks have 1%, return on assets; calculate the return on equity is ROE = ROA*Equity multiplier, so:

ROE(X) = 0.01*10 = 0.1,

ROE(Y) = 0.01*25 = 0.25.